a compound Y with chemical formula as shown below:


CH2=CHCOOCH3


b) calculate the percentage composition of carbon in the compound Y.


1
Expert's answer
2021-08-09T07:49:53-0400

we will find it by the formula:

"w(C)=\frac{т\times Ar(C)}{Mr(CH2 = CHCOOCH3)}=\frac{4\times12}{86.09}=0.56" or 56%

Mr(CH2 = CHCOOCH3)=86.09